# Detecting Shearless Phase-Space Transport Barriers in Global Gyrokinetic Turbulence Simulations with Test Particle Map Models

### v1

@@ 4.2 Shearless Barrier Reconnection
Finally, using these silhouettes, we can track if the topological features of the ion gyrocenter distribution function $F_{i}$ correlate with the avalanches observed in the physical fields. This is shown in the bottom panels (c) and (d) of Figure 10, where the silhouettes are plotted in a color panel. $P_{blob}$ and $P_{hole}$ correspond to the persistence of local maxima and minima of $F_{i}$. Stronger persistence is shown with brighter colors, with the x-axis corresponding to the time and the y-axis approximately corresponding to radial location.
@@
The phase space blobs show a clear correlation with the incoming avalanches from the edge, showing that these avalanches are tied to local topological features of the ion gyrocenter distribution function. These blobs consistently terminate at a particular value of $F_{i}\approx 2.75$, which corresponds to the value of $F_{i}$ aligned with the shearless torus. This indicates that the shearless phase-space barrier reconnection event observed in panel (a) is a generic feature of inward-propagating avalanches. Note that there are faint bands of outward propagating phase space holes visible in panel (d), although these are not as pronounced as the inward propagating phase space blobs. This is consistent with the expectation that turbulence spreading would be more relevant to the larger amplitude fluctuations coming from the edge.
@@
We conclude this section with a brief speculation about the potential mechanism by which shearless transport barriers could lead to the arrest of turbulence propagation. In Pratt & Stern (1986); Pratt (1988) it was proposed that eddy detachment in the Gulf Stream could be understood through the idealized model following the contour dynamics of a single potential vorticity (PV) front. The key process that leads to eddy detachment in the model is vortex induction due to the north-south displacement of the PV front. The strong PV jump across the jet correspondingly leads to a strong nonlinear vortex flow induced by displacements of the jet, leading to wave breaking and subsequent eddy detachment.
@@
In magnetized plasmas, the PV can be identified with the ion gyrocenter density (McDevitt et al., 2010; Gürcan & Diamond, 2015). Thus, the situation observed in the XGC simulations of a sudden increase in the ion gyrocenter density gradients across the zonal $E\times B$ jet is closely analogous to the situation of the PV jump across the oceanic jet. Physically, displacements of the plasma across the $E\times B$ jet would correspond to a sudden intrusion of lower or higher ion gyrocenter density across the shearless region. Since the adiabatic parallel electron response is relevant for the outer core, the expectation would be that this sudden density change would induce significant blob spin (Myra et al., 2004), i.e. vortical $E\times B$ flow about the blob (or hole) center. Spin mitigates curvature-induced charge polarization in blobs, which is precisely the process responsible for driving the ITG instability, suggesting a potential physical mechanism for the arrest of turbulence propagation at the shearless region. For future work, it may be interesting to use blob seeding experiments in simulations (Cheng et al., 2023) to perform controlled numerical experiments to more precisely determine the ways in which shearless regions impact the propagation of turbulence.
@@ 5.1 Applicability to Collisional and Broadband Turbulence
We first discuss how the reliance of shearless transport barriers on collisionless test particle dynamics used in the model can be relaxed. Shearless transport barriers were originally studied in the fluid limit of drift wave turbulence, whereas the collisionless limit studied here focused on trapped particle dynamics which are completely absent from the fluid limit. Thus, this suggests that shearless transport barriers could play different roles throughout a range of collisionalities in plasma turbulence. To identify shearless transport barriers in regimes with higher collisionality, we highlight two theoretical approaches. The first is the relationship between the local Cauchy-Green tensor, which integrates the total local deformation of a Lagrangian parcel, and diffusion. In fluid models, a certain minimum deformation principle can be used to define a variational notion of shearless material transport barrier (Farazmand et al., 2014), across which the combined effect of diffusion and stretching due to advection will be minimum. Another approach is to use stochastic Lagrangian trajectories, in which the effect of both collisions and advective transport is accounted for through a stochastic perturbation to the test particle dynamics. In studies including Kwon et al. (2000); Cao & Qi (2024), it is noted that shearless tori can continue to act as effective transport barriers in the presence of collisions.
@@
The first remark is on the relationship between the single-mode assumption and particle stochasticity. Particle stochasticity is a major pillar of quasilinear theory (Diamond et al., 2010), and typically a broad spectrum of modes is invoked in order to satisfy the resonance overlap criterion necessary for particle stochasticity. For the toroidal mode number $n=39$ of the ballooning mode studied here, rational surfaces are closely spaced. The representation (5) keeps a broad range of poloidal harmonics active, and evidently from the Poincaré sections in figure 6 the island overlap criterion is satisfied over most of the region of interest. Thus, even a single tertiary mode is able to drive significant transport by itself, potentially relaxing driving gradients and temporarily “outcompeting” other tertiary modes in the immediate vicinity. Since wave-kinetic models would tend to predict the localization of drift-wave modes around minima and maxima of the $E\times B$ rotation rate, this could provide a generic mechanism by which energy is accumulated into a small number of tertiary modes, enabling the usage of a single- or few-mode model. For example, Cao & Qi (2024) demonstrates how chaotic mixing could reinforce the amplitude of vortical flows associated with tertiary modes in a fluid slab model of resistive drift-wave turbulence.
@@
As discussed in §3.2, González-Enríquez et al. (2014) develops a systematic classification for shearless invariant tori in higher dimension using singularity theory. In their classification, non-degenerate shearless tori in planar maps correspond to the “fold” singularity, originating from the fact that the lowest-order Birkhoff normal form of the Hamiltonian is cubic in the action variable in the vicinity of a non-degenerate shearless torus. As discussed in 3.2, shearless tori in higher dimension occur when the local shear $D\omega(J)$ is singular, or equivalently when $\operatorname{det}(D\omega(J))=0$. Intuitively, one can expect that the zero level set of the scalar function $\operatorname{det}(D\omega(J))$ over the action variables $J$ as well as any parameters $\lambda$ in the system forms a co-dimension 1 manifold, which is indeed the case for the fold singularity.
Since invariant tori are usually $N$-dimensional for $2N$-dimensional Hamiltonian systems, there is normally no expectation that surviving KAM tori for the usual twist case will act as global barriers to transport. Intuitively, it is easy for paths in $2N$-dimensional phase space to avoid $N$-dimensional barriers. The importance of the fold singularity is that it implies that shearless tori in higher dimension are organized into co-dimension 1 families. Supposing that invariant tori in this co-dimension 1 family are somehow “robust” to perturbations like their planar map cousins, then surviving shearless KAM tori could act as a global barrier to transport. This may explain why the severe truncation of the model electrostatic perturbations in the simulations to a single mode can capture qualitative features of $F_{i}$ in figure 9; the addition of other modes with sufficiently small amplitude may simply perturb the family of shearless phase-space transport barriers observed in the test particle model over different values of $v_{\parallel},v_{\perp}$ without destroying them. Given advances in computational methods for detecting invariant tori in higher-dimensional systems, see e.g. Haro et al. (2016); Ruth et al. (2025), it may be within reach to explore the robustness of such structures, which may be a fertile avenue for future work.
@@ 5.2 Extension to Other Turbulence Regimes
Moving on to pedestal turbulence, a wide array of ion-scale modes can be destabilized including microtearing modes (MTMs), ITGs, trapped electron modes (TEMs), kinetic ballooning modes (KBMs), and drift-Alfvén waves (DAWs) (Kotschenreuther et al., 2019; Diallo & Laggner, 2020). Moreover, there are several scenarios where pedestal-localized modes are observed to organize into weakly or quasi-coherent bands, the latter sometimes being referred to as “washboard modes” if multiple bands are visible. Fluctuations are also often observed to radially localize somewhere close to strong gradient region, which again corresponds to the shearless region. MTMs in H-mode (Hatch et al., 2016, 2021) have been observed to occur as washboard modes, with their radial localization determined by a well in the electron diamagnetic frequency. The weakly coherent mode (WCM) in I-mode (Liu et al., 2016; Herschel et al., 2024), although its nature and role in regulating transport in I-mode is still unclear, is also a weakly coherent electromagnetic mode which is localized to the pedestal in the presence of a temperature barrier. In quiescent H-modes (QH-mode) and wide pedestal quiescent H-modes (WPQH-mode), fluctuation measurements also detect fluctuations localized to the $E_{r}$ well and linear analyses have shown the existence of modes which match this degree of radial localization (Li et al., 2022).
@@ 6 Summary
We demonstrated how a map model of the collisionless gyrokinetic test particle dynamics was able to identify shearless tori that corresponded with this shearless transport barrier. This model was constructed by extracting the dominant fluctuation mode in the shearless region, which was an intermediate-$n$ ($n=39$) electrostatic ballooning mode with a finite radial envelope localized to the zonal jet. Non-monotonic behavior in the kinetic safety factor profile, $q_{kin}$, was identified as a necessary ingredient for shearless transport barriers to exist. In particular, under the usual conditions of magnetic shear, this condition is only satisfied for the trapped particle portion of phase space. Poincaré sections were used to visually identify the presence of shearless invariant tori in the model, and direct signatures of shearless tori were identified in the ion gyrocenter distribution function $F_{i}$ taken from XGC. A transmissivity analysis showed that the net effect of the shearless tori was to reduce radial particle transport by a factor of $\sim 2$ in the trapped portions of phase space, potentially explaining the observed increase in ion gyrocenter density gradients. Finally, the arrest of turbulence propagation was linked to phase-space barrier reconnection events in $F_{i}$, and “blob spin” was proposed as the physical mechanism for this arrest.
@@ Acknowledgements
The authors thank D.R. Hatch and P.J. Morrison for fruitful comments on this work. This material is based upon work supported by the US Department of Energy, Office of Science, Contract Numbers DE-FG02-04ER54742 and DE-AC02-09CH11466. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award FES-ERCAP002927.
@@
The code associated with this work can be found at https://github.com/Maplenormandy/c1lgkt

### v2

@@ 4.2 Shearless Barrier Breakdown via Eddy Detachment
Using these silhouettes, we can track how topological features of the ion gyrocenter distribution function $F_{i}$ correlate with the avalanches observed in the physical fields. We focus on the detection of local maxima and minima of the ion gyrocenter distribution function $F_{i}$ in each of the 16 computational toroidal planes over a range of values of time $t$. When restricted to $E_{\mu,K}$, this corresponds to finding local maxima and minima of planar functions. We use the simplex tree and representations modules of the GUDHI library (Boissonnat and Maria, 2014; Maria, 2025; Dlotko, 2025) to compute the persistence diagrams and silhouettes. The results of this analysis is shown in panels (c) and (d) of Figure 10, where the silhouettes are plotted in a color panel. $P_{blob}$ and $P_{hole}$ correspond to the persistence of local maxima and minima of $F_{i}$, and stronger persistence is shown with brighter colors. Following the earlier discussion, peaks in the persistence along the vertical axis of these plots correlate with radial location of bands of maxima/minima in $F_{i}$, while evolution of the peaks in the horizontal axis give the time evolution of these radial bands.
@@
Focusing on panel (c), the phase space blobs show clear correlation with incoming avalanches from the edge, showing that these avalanches are tied to local topological features of the ion gyrocenter distribution function. These blobs consistently terminate at a particular value of $F_{i}\approx 2.75$, which corresponds to the value of $F_{i}$ aligned with the shearless torus. This indicates that the shearless phase-space barrier eddy detachment event observed in panel (a) is a generic feature of inward-propagating avalanches. Note that panel (d) shows faint bands of outward propagating phase space holes, although these are not as pronounced as the inward propagating phase space blobs. This is consistent with the expectation that turbulence spreading would be more relevant to the larger amplitude fluctuations coming from the edge.
@@
We conclude this section with a brief speculation about the potential mechanism by which shearless transport barriers could lead to the arrest of turbulence propagation. In Pratt and Stern (1986); Pratt (1988) it was proposed that eddy detachment in the Gulf Stream could be understood through the idealized model following the contour dynamics of a single potential vorticity (PV) front. The key process that leads to eddy detachment in the model is vortex induction due to the north-south displacement of the PV front. The strong PV jump across the jet correspondingly leads to a strong nonlinear vortex flow induced by displacements of the jet, leading to wave breaking and subsequent eddy detachment. We remark that this can also be interpreted as a type of reconnection for meandering tori, potentially related to the breakup of shearless tori in dynamical systems models (Wurm et al., 2005).
@@
In magnetized plasmas, the PV can be identified with the ion gyrocenter density (McDevitt et al., 2010; Gürcan and Diamond, 2015). Thus, the situation observed in the XGC simulations, namely a sudden increase in the ion gyrocenter density gradient across the zonal $E\times B$ jet, is closely analogous to the situation of the PV jump across the oceanic jet. Physically, displacements of the plasma across the $E\times B$ jet would correspond to a sudden intrusion of lower or higher ion gyrocenter density across the shearless region. Since the adiabatic parallel electron response is relevant for the outer core, the expectation would be that this sudden density change would induce significant blob spin (Myra et al., 2004), i.e. vortical $E\times B$ ‘eddy’ flow about the blob (or hole) center. Spin mitigates curvature-induced charge polarization in blobs, which is precisely the process responsible for driving the ITG instability, suggesting a potential physical mechanism for the arrest of turbulence propagation at the shearless region. For future work, it may be interesting to use blob seeding experiments in simulations (Cheng et al., 2023) to perform controlled numerical experiments to more precisely determine the ways in which shearless regions impact the propagation of turbulence.
@@ 5.1 Applicability to Collisional and Broadband Turbulence
We first discuss how the reliance of shearless transport barriers on collisionless test particle dynamics used in the model can be relaxed. Shearless transport barriers were originally studied in the fluid limit of drift wave turbulence, whereas the collisionless limit studied here focused on trapped particle dynamics which are completely absent from the fluid limit. Thus, this suggests that shearless transport barriers could play different roles throughout a range of collisionalities in plasma turbulence. To identify shearless transport barriers in regimes with higher collisionality, we highlight two theoretical approaches. The first is the relationship between the local Cauchy-Green tensor, which integrates the total local deformation of a Lagrangian parcel, and diffusion. In fluid models, a certain minimum deformation principle can be used to define a variational notion of shearless material transport barrier (Farazmand et al., 2014), across which the combined effect of diffusion and stretching due to advection will be minimum. Another approach is to use stochastic Lagrangian trajectories, in which the effect of both collisions and advective transport is accounted for through a stochastic perturbation to the test particle dynamics. In studies including Kwon et al. (2000); Cao and Qi (2024), it is noted that shearless tori can continue to act as effective transport barriers in the presence of collisions.
@@
The first remark is on the relationship between the single-mode assumption and particle stochasticity. Particle stochasticity is a major pillar of quasilinear theory (Diamond et al., 2010), and typically a broad spectrum of modes is invoked in order to satisfy the resonance overlap criterion necessary for particle stochasticity. For the toroidal mode number $n=39$ of the ballooning mode studied here, rational surfaces are closely spaced. The representation (5) keeps a broad range of poloidal harmonics active, and evidently from the Poincaré sections in figure 6 the island overlap criterion is satisfied over most of the region of interest. Thus, even a single tertiary mode is able to drive significant transport by itself, potentially relaxing driving gradients and temporarily “outcompeting” other tertiary modes in the immediate vicinity. Since wave-kinetic models would tend to predict the localization of drift-wave modes around minima and maxima of the $E\times B$ rotation rate, this could provide a generic mechanism by which energy is accumulated into a small number of tertiary modes, enabling the usage of a single- or few-mode model. For example, Cao and Qi (2024) demonstrates how chaotic mixing could reinforce the amplitude of vortical flows associated with tertiary modes in a fluid slab model of resistive drift-wave turbulence.
@@
As discussed in §3.2, González-Enríquez et al. (2014) develops a systematic classification for shearless invariant tori in higher dimension using singularity theory. In their classification, non-degenerate shearless tori in planar maps correspond to the “fold” singularity, originating from the fact that the lowest-order Birkhoff normal form of the Hamiltonian is cubic in the action variable in the vicinity of a non-degenerate shearless torus. In higher dimension, shearless tori occur when the local shear $D\omega(J)$ is singular, or equivalently when $\operatorname{det}(D\omega(J))=0$. Intuitively, one can expect that the zero level set of the scalar function $\operatorname{det}(D\omega(J))$ over the action variables $J$ as well as any parameters $\lambda$ in the system forms a co-dimension 1 manifold, which is indeed the case for the fold singularity.
Since invariant tori are usually $N$-dimensional for $2N$-dimensional Hamiltonian systems, there is normally no expectation that surviving KAM tori for the usual twist case will act as global barriers to transport. Intuitively, it is easy for paths in $2N$-dimensional phase space to avoid $N$-dimensional barriers when $N>1$. The importance of the fold singularity is that it implies that shearless tori in higher dimension are organized into co-dimension 1 families. Conjecturing that higher-dimensional ($N>1$) invariant tori in co-dimension 1 fold families are “robust” to perturbations like their planar map ($N=1$) cousins, then surviving shearless KAM tori could act as global barriers to transport. In collisionless gyrokinetics, where the gyrophase is an ignorable coordinate and the magnetic moment $\mu$ can be taken as a fixed parameter of the particle dynamics, the $N=2$ case would be relevant.
We also conjecture that the persistence of fold tori may be related to the ability of the model electrostatic perturbations to capture qualitative features of $F_{i}$ in figure 9. Although the single mode approximation represents a severe truncation of the fluctuation spectrum observed in the simulations, the addition of other modes with sufficiently small amplitude may simply perturb the family of shearless phase-space transport barriers observed in the test particle model over different values of $v_{\parallel},v_{\perp}$ without destroying them. Given advances in computational methods for detecting invariant tori in higher-dimensional systems, see e.g. Haro et al. (2016); González et al. (2022); Ruth et al. (2025), it may be within reach to explore the robustness of such structures, which may be a fertile avenue for future work.
@@ 5.2 Extension to Other Turbulence Regimes
Moving on to pedestal turbulence, a wide array of ion-scale modes can be destabilized including microtearing modes (MTMs), ITGs, trapped electron modes (TEMs), kinetic ballooning modes (KBMs), and drift-Alfvén waves (DAWs) (Kotschenreuther et al., 2019; Diallo and Laggner, 2020). Moreover, there are several scenarios where pedestal-localized modes are observed to organize into weakly or quasi-coherent bands, the latter sometimes being referred to as “washboard modes” if multiple bands are visible. Fluctuations are also often observed to radially localize somewhere close to strong gradient region, which again corresponds to the shearless region. MTMs in H-mode (Hatch et al., 2016, 2021) have been observed to occur as washboard modes, with their radial localization determined by a well in the electron diamagnetic frequency. The weakly coherent mode (WCM) in I-mode (Liu et al., 2016; Herschel et al., 2024), although its nature and role in regulating transport in I-mode is still unclear, is also a weakly coherent electromagnetic mode which is localized to the pedestal in the presence of a temperature barrier. In quiescent H-modes (QH-mode) and wide pedestal quiescent H-modes (WPQH-mode), fluctuation measurements also detect fluctuations localized to the $E_{r}$ well and linear analyses have shown the existence of modes which match this degree of radial localization (Li et al., 2022).
@@ 6 Summary
We demonstrated how a map model of the collisionless gyrokinetic test particle dynamics was able to identify shearless tori that corresponded with this shearless transport barrier. This model was constructed by extracting the dominant fluctuation mode in the shearless region, which was an intermediate-$n$ ($n=39$) electrostatic ballooning mode with a finite radial envelope localized to the zonal jet. Non-monotonic behavior in the kinetic safety factor profile, $q_{kin}$, was identified as a necessary ingredient for shearless transport barriers to exist. In particular, under the usual conditions of magnetic shear, this condition is only satisfied for the trapped particle portion of phase space. Poincaré sections were used to visually identify the presence of shearless invariant tori in the model, and direct signatures of shearless tori were identified in the ion gyrocenter distribution function $F_{i}$ taken from XGC. A transmissivity analysis showed that the net effect of the shearless tori was to reduce radial particle transport by a factor of $\sim 2$ in the trapped portions of phase space, potentially explaining the observed increase in ion gyrocenter density gradients. Finally, the arrest of turbulence propagation was linked to phase-space barrier eddy detachment events in $F_{i}$, and “blob spin” was proposed as the physical mechanism for this arrest.
@@ Acknowledgements
The authors thank D.R. Hatch and P.J. Morrison for fruitful comments on this work. Several anonymous individuals are also thanked for their feedback. This work was supported by the US Department of Energy, Office of Science (N.M.C. contract number DE-FG02-04ER54742), (H.Z. and T.S. contract number DE-AC02-09CH11466); and the São Paulo Research Foundation (FAPESP), Brasil (G.C.G. Process Number #2024/02591-0). This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award FES-ERCAP002927.
Declaration of Interests
The authors report no conflict of interest.
@@
The code that support the findings of this study are openly available in https://github.com/Maplenormandy/c1lgkt